OpenStudy (anonymous):

How do I solve this rational expression? (x/4x-8)-(3x+1/3x-6)

OpenStudy (anonymous):



OpenStudy (anonymous):

this is the question ?

OpenStudy (lukecrayonz):

(x/(4x-8))-(3x+1)/(3x-6)

OpenStudy (anonymous):

yes @Coolsector

OpenStudy (anonymous):

so common denominator: [x(3x-6) - (3x+1)(4x-8)]/[(3x-6)(4x-8)] = (3x^2-6x - 12x^2-4x+24x+8) / [(3x-6)(4x-8)] = (-9x^2+14x+8) / [(3x-6)(4x-8)] = (-9x^2 + 18x - 4x + 8) / [(3x-6)(4x-8)] = (-9x(x+2)-4(x-2)) / [(3x-6)(4x-8)] = (-9x-4)(x-2)/[(3x-6)(4x-8)] = (-9x-4)(x-2) / [4(3x-6)(x-2)] = -(9x+4) /[12(x-2)] and x is not 2.

OpenStudy (isaiah.feynman):

@Katelynmorris here is a general procedure for solving rational expressions. \[\frac{ a }{ b }-\frac{ c }{ d } = \frac{ ad-bc }{ bd }\]

OpenStudy (isaiah.feynman):

Did you solve it yet? @Katelynmorris

OpenStudy (anonymous):

Not yet.

OpenStudy (isaiah.feynman):

Okay.